The Country Faire Company wants achieve $5 million in additional sales. They believe that they can raise $300,000 in additional funding from investors and that vendors will extend an additional $200,000 in credit. If the firm pays out 40% of earnings as dividends and profit margin is 10% on the additional sales, then how much can they afford to spend on assets to make this plan work?

$375,000
$500,000
$600,000
$675,000
$800,000
Answer #1

Net income=sales*profit margin

=5000000*10%

=500000

retained earnings=net income*(1-dividend rate)=500000*(1-40%)=300000

how much can they afford to spend on assets to make this plan work

=raise additional funding from investors+retained earnings+credit from vendors

=300000+300000+200000

=800000

the above is answer..